Need help to crate a word math problem not sure how to do?
If a daily cost of breakfast is $10.00 and Lunch is $22.00 and Dinner is $35.00 the sum = X
(b+l+d)= X
Now if I budgeted B=$87.00 then how do I write out the loss?
X/b= ??
I know this is all wrong HELP!!!

Answer by JulietG(1812)

B + L + D = total amount spent
PD (per diem) = $87
.
10 + 22 + 35 = 67
If you budgeted 87, you don't have a loss. You have an excess of $20
.
PD-(B+L+D) = profit/loss
.
If, however, you spent 65 on dinner, then you would have 97 total.
87-97 is -10, which is generally written in accounting as <10>.